# Real-time Verification and Refinement of Language Model Text Generation

### v4

@@ 5 Conclusion
In this paper, we introduce Streaming-VR, a novel approach aimed at improving the accuracy and efficiency in language model text generation. Unlike traditional methods solely relying on the final response, Streaming-VR performs real-time verification and correction of erroneous token sequences as they are being produced, with external models simultaneously with answer generation. This prevents error propagation in the early stage and reduces the errors at the end by minimizing the likelihood of compounding inaccuracies, then significantly enhances the efficiency of answer refinement. Extensive experiments for two different QA datasets have clearly demonstrated that Streaming-VR consistently achieves remarkably higher efficiency without compromising response quality.
@@
Despite the improvements introduced by our method, Streaming-VR, which enhances both the efficiency and effectiveness of verification and refinement in language model text generation by intervening during intermediate answer generation, there remain promising opportunities for enhancing the answer verifier. Specifically, the primary challenge is the lack of dedicated datasets for answer verification, particularly those suited for real-time scenarios. To address this, we automatically augmented data by paraphrasing sentences or introducing errors by an LLM. However, while effective, this approach carries the risk of mislabeling. Therefore, future work could focus on developing new datasets that are carefully annotated with a diverse range of answers ensuring more accurate verification and reducing the risk of incorrect labeling. Additionally, we can further extend these datasets to include fine-grained labels for multiple classes, rather than just binary ones, to accommodate different types of errors and apply adaptive strategies for subsequent refinement after verification.
@@ Appendix A Implementation Details
In our experiments, we employ two open-source LLMs Mistral 7B (Jiang et al., 2023a) and LLaMA-3.1 8B (Dubey et al., 2024) via Hugging Face (Wolf et al., 2020) API and GPT-4o (Achiam et al., 2023) which is accessible via OpenAI API, representing relatively small, medium, and large models, respectively. Here, these models are never fine-tuned or further trained except for their roles in verification. For the overall Streaming-VR pipeline, LLaMA-3.1 8B functions as the primary backbone language model to generate answers for given queries, while all three models are employed for verification or refinement for experiments.
@@
We fine-tuned Mistral 7B and LLaMA-3.1 8B as verifiers using augmented training data derived from the ASQA and QuoteSum datasets. Each verifier was trained for five epochs on its respective training set, with a learning rate of 1e-5 and the AdamW (Loshchilov and Hutter, 2019) optimizer.
To generate augmented data for false-labeled sentences, we embedded fake information into true sentences using GPT-4o, adjusting the temperature to , , and to create diverse forms of inaccuracies. Rather than synthesizing entirely new sentences with large language models, which risk introducing unrelated hallucinations, we adopted this targeted augmentation strategy as a more reliable approach. This method proved highly effective in training verifiers to identify hallucinations, delivering exceptional results that highlight the importance of Streaming-VR in improving efficiency while preserving answer quality. The specific prompt used to generate incorrect information for each sentence (Sentence) in the provided answer (Answer) to a given question (Question) is detailed below.
The final test results of the finetuned verifiers used in the experiments, including a Random baseline that selects verification results arbitrarily, are presented in Table 4. For the entire pipeline, we establish a constraint that prohibits the use of any other verifier models larger than the answer generation model. This decision is based on the principle that the verifier should not exceed the capabilities of the response model, as the verifier serves merely as a supplementary tool for identifying mistakes. This reflects our considerations regarding computational costs and efficiency.
@@ Appendix B Additional Experimental Results
In Table 5, we present additional results for other baselines on ASQA, Self-RAG (Asai et al., 2024), one of the most representative methods with their trained models publicly available. Self-RAG performs on-demand retrieval of external information via a specialized retrieval token, followed by a critique of the generated output to refine it. When we compare the baseline results with the similar model size, Mistral 7B demonstrates significantly superior performance to Self-RAG even without the help of refinement.
@@ Appendix C Time-Efficiency of Streaming-VR
Here, we assume that the inference time of verification model per call, , is the same regardless of the input length for simple comparison, and denotes the token generation time of the refiner. Since the verifier does not generate tokens, it follows that . Furthermore, as shown in Table 2, with an average of seven sentences per answer , we observe that . Consequently, we can conclude that due to the following inequality:
@@
Appendix D Case Study
@@
We have conducted case studies to examine the advantages of Streaming-VR in greater detail. In Table 6, we present the results of various methods for the query: "Who does sam marry in foyle’s war?", a British detective drama television series broadcast on ITV. The answer to the question is Adam Wainwright. However, this question introduces ambiguity because the answer could refer either to Max Brown, who portrayed Adam Wainwright in Season 6, or Daniel Weyman, who was recast as Adam Wainwright in Seasons 7 and 8.
As shown in Table 6, the original answer generated by the naïve LLaMA-3.1 8B model reflects the model’s difficulty in understanding due to grammatical errors and ambiguous information. However, when using the methods that enable the refinement through the verification (with red to be False and blue to be True), Full-VR still contains an error regarding the name of the broadcasting center, whereas Streaming-VR successfully corrects every error in the answer. Also, we observe that the answer generated and refined by Streaming-VR preserves the structure of the original answer, whereas Full-VR generates an entirely new response, significantly altering the sentence structure.

### v5

@@ 5 Conclusion
In this paper, we introduced Streaming-VR, a novel approach aimed at improving the accuracy and efficiency in language model text generation. Unlike traditional methods solely relying on the final response, Streaming-VR performs real-time verification and correction of erroneous token sequences as they are being produced, with external models simultaneously with answer generation. This prevents error propagation in the early stage and reduces the errors at the end by minimizing the likelihood of compounding inaccuracies, then significantly enhances the efficiency of answer refinement. Extensive experiments for two different question-answering datasets have clearly demonstrated that Streaming-VR consistently achieves remarkably higher efficiency without compromising accuracy.
@@
Despite the improvements from Streaming-VR, which enhances both the efficiency and effectiveness of verification and refinement in language model text generation by intervening during intermediate answer generation, there remain promising opportunities for enhancing the answer verifier. Specifically, the primary challenge is the lack of dedicated datasets for answer verification, particularly those suited for real-time scenarios. To address this, we automatically augmented data by paraphrasing sentences or introducing errors by an LLM. However, while effective, this approach carries the risk of mislabeling. Therefore, future work could focus on developing new datasets that are carefully annotated with a diverse range of answers ensuring more accurate verification and reducing the risk of incorrect labeling. Additionally, we can further extend these datasets to include fine-grained labels for multiple classes, rather than just binary ones, to accommodate different types of errors and apply adaptive strategies for subsequent refinement after verification.
@@
Acknolwedgments
This work was supported by the Institute for Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (RS-2019-II190075, Artificial Intelligence Graduate School Program (KAIST), and RS-2022-II220713, Meta-learning Applicable to Real-world Problems), the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (RS-2023-00256259), the grant of the Korea Machine Learning Ledger Orchestration for Drug Discovery Project (K-MELLODDY) funded by the Ministry of Health & Welfare and the Ministry of Science and ICT, Republic of Korea (RS2024-00460870), Institute of Information & Communications Technology Planning & Evaluation (IITP) with the grant funded by the Ministry of Science and ICT (MSIT) of the Republic of Korea in connection with the Global AI Frontier Lab International Collaborative Research (RS-2024-00469482 & RS-2024-00509279), the Artificial Intelligence Industrial Convergence Cluster Development Project funded by the Ministry of Science and ICT (MSIT, Korea) & Gwangju Metropolitan City, and i-Scream Media.
@@ Appendix A Implementation Details
In our experiments, we employ two open-source LLMs Mistral 7B (Jiang et al., 2023a) and LLaMA-3.1 8B (Dubey et al., 2024) via Hugging Face Hub (Wolf et al., 2020) and GPT-4o (Achiam et al., 2023) which is accessible via OpenAI API, representing relatively small, medium, and large models, respectively. Here, these models are never fine-tuned or further trained except for their roles in verification. For the overall Streaming-VR pipeline, LLaMA-3.1 8B functions as the primary backbone language model to generate answers for given queries, while all three models are employed for verification or refinement for experiments.
@@
We fine-tuned Mistral 7B and LLaMA-3.1 8B as verifiers using augmented training data from ASQA and QuoteSum datasets. Each verifier was trained for five epochs on its respective training set, with a learning rate of 1e-5 and the AdamW (Loshchilov and Hutter, 2019) optimizer. To generate augmented data for false-labeled sentences, we embedded fake information into true sentences using GPT-4o, adjusting the temperature to , , and to create diverse forms of inaccuracies. Rather than synthesizing entirely new sentences with large language models, which risk introducing unrelated hallucinations, we adopted this targeted augmentation strategy as a more reliable approach. This method proved highly effective in training verifiers to identify hallucinations, delivering exceptional results that highlight the importance of Streaming-VR in improving efficiency while preserving answer quality. The specific prompt used to generate incorrect information for each sentence (Sentence) in the provided answer (Answer) to a given question (Question) is detailed in the prompt box as follows.
@@
The final test results of the finetuned verifiers used in the experiments, including a Random baseline that selects verification results arbitrarily, are presented in Table˜4. For the entire pipeline, we establish a constraint that prohibits the use of any other verifier models larger than the answer generation model. This decision is based on the principle that the verifier should not exceed the capabilities of the response model, as the verifier serves merely as a supplementary tool for identifying mistakes. This reflects our considerations regarding computational costs and efficiency.
@@ Appendix B Additional Experimental Results
Additional Baselines
In Table˜5, we present additional results for other baselines on ASQA, Self-RAG (Asai et al., 2024), one of the most representative methods with their trained models publicly available. Self-RAG performs on-demand retrieval of external information via a specialized retrieval token, followed by a critique of the generated output to refine it. When we compare the baseline results with the similar model size, Mistral 7B demonstrates significantly superior performance to Self-RAG even without the help of refinement.
@@ Appendix C Time-Efficiency of Streaming-VR
Here, we assume that the inference time of the verification model per call, , is the same regardless of the input length for simple comparison, and denotes the token generation time per token of the refiner. Since the verifier does not generate tokens, it follows that . Furthermore, as shown in Table˜2, with an average of seven sentences per answer , we observe that . Consequently, we can conclude that due to the following inequality:
@@
Appendix D GPU memory
In a practical deployment setting, Full-VR and Streaming-VR exhibit the same maximum GPU memory usage. To be specific, as the maximum number of tokens that Full-VR and Streaming-VR should handle is exactly the same, which depends on the generated (and refined) answer from the original response model. Quantitatively, when using LLaMA-3.1 8B for answer generation and Mistral 7B for verification, it requires approximately 43,900 (16,035 + 27,865) MB of GPU memory.
Appendix E Case Study
@@
We have conducted case studies to examine the advantages of Streaming-VR in greater detail. In Table˜6, we present the results of various methods for the query: "Who does sam marry in foyle’s war?", a British detective drama television series broadcast on ITV. The answer to the question is Adam Wainwright. However, this question introduces ambiguity because the answer could refer either to Max Brown, who portrayed Adam Wainwright in Season 6, or Daniel Weyman, who was recast as Adam Wainwright in Seasons 7 and 8.
As shown in Table˜6, the original answer generated by the naïve LLaMA-3.1 8B model reflects the model’s difficulty in understanding due to grammatical errors and ambiguous information. However, when using the methods that enable the refinement through the verification (with red to be False and blue to be True), Full-VR still contains an error regarding the name of the broadcasting center, whereas Streaming-VR successfully corrects every error in the answer. Also, we observe that the answer generated and refined by Streaming-VR preserves the structure of the original answer, whereas Full-VR generates an entirely new response, significantly altering the sentence structure.